a) 7 cm (b) 14 cm (c) 3.5 cm d) 17.5 cm

a) 7 cm
(b) 14 cm
(c) 3.5 cm
d) 17.5 cm

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Areas Related To Circles Class 10 MCQ With Answer: dExplaination: Reason: Area of circle = 154 cm² ⇒ nr² = 154 cm2 ⇒ 227 × r² = 154 ⇒ r² = 154 × 227 ⇒ r² = 7 × 7 = 49 ∴ r = √49 = 7